# Fair Graph Machine Learning under Adversarial Missingness Processes

### v4

@@ Appendix A Technical Appendices and Supplementary Material
$\min_{\theta_{C}}\sup_{\hat{s}}|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$
Let us assume that there can be $m$ different $\hat{s}$ that produce sensitive attributes $\hat{s}_{1},\hat{s}_{2},...,\hat{s}_{m}$. If we let $D_{i}$ be the uniform distribution over each $s_{i}$, then BFtS minimizes the $|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$ in the worst case over these different distributions $D_{i}$. Therefore, $f^{*}_{C}$ is a minimax estimator. The objective of Eq. 7 with optimal $f^{*}A,f^{*}I$ and $f^{*}_{C}$ becomes:
@@
Therefore, $f_{C}$ is a minimax estimator. As $f_{C}$ is a minimax estimator, the maximal $\Delta DP$ of BFtS is minimum amongst all estimators of $s$
At the global minima of $f_{C}$, $|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$ will be minimum. Its minimum value is 0. Therefore, $p(\mathbf{h}|\hat{s}=1)-p(\mathbf{h}|\hat{s}=0)=0$. The optimal $f^{*}_{C}$ achieves demographic parity for the worst case.
@@
If $s^{\prime}$ is a separate imputation independent of $f_{C}$ and $f_{A}$, then the min max game according to [14] is the following:
$\begin{split}\min_{\theta_{C}}\max_{\theta_{A}}~&\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|s^{\prime}=1)}[\log f_{A}(\mathbf{h})]\\ &+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|s^{\prime}=0)}[\log(1-f_{A}(\mathbf{h}))].\end{split}$
With optimal $f^{*}_{A}$, the optimization problem simplifies to-
$\min_{\theta_{C}}-\log 4+JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0)).$
With independent imputation $s^{\prime}$, the adversary $f_{A}$ tries to approximate the lower bound of Jensen Shannon Divergence [71]. In BFtS, the adversarial imputation $f_{I}$ tries to approximate the upper bound of the JS divergence, and the classifier $f_{C}$ tries to minimize the upper bound on the JS divergence provided by the adversarial imputation (see Proposition 1). Therefore,
@@
Let us consider a scenario when the independent missing sensitive value imputation [14] fails to converge. There are three possible reasons for the convergence to fail:
@@
In scenario $1$ and $2$ the supports of $p(\mathbf{h}|s^{\prime}=1)$ and $p(\mathbf{h}|s^{\prime}=0)$ are disjoint. According to [14], the optimization for the classifier with an optimal adversary is
$\min_{\theta_{C}}-\log 4+2JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$
@@
In scenario 3, $\hat{s}$ is uniform. Therefore, $p(\mathbf{h}|s^{\prime}=1)$ will be equal to $p(\mathbf{h}|s^{\prime}=0)$ which results in $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ being equal to 0. Moreover, the classifier may assign all training samples to a single class and yet have an adversary that may not be able to distinguish the sensitive attributes of the samples as $\hat{s}$ is random. If the classifier assigns all samples to the majority class, it will achieve low classification loss $L_{C}$ along with minimum $L_{A}$. The training of the classifier may get stuck in this local minima (see Eq. 4). This phenomenon is similar to the mode collapse of GANs [65].
@@
Algorithm 1 is a high-level description of the key steps applied for training BFtS. It receives the graph $\mathcal{G}$, labels $y$, labeled nodes $\mathcal{V}_{L}$, nodes with observed sensitive attributes $\mathcal{V}_{S}$ and hyperparameters, $\alpha$ and $\beta$ as inputs and outputs the GNN classifier $f_{C}$, sensitive attribute predictor, $f_{A}$ and missing data imputation GNN $f_{I}$. We first estimate $\hat{si}$ and then fix $\theta_{C}$ and $\theta_{A}$ to update $\theta_{I}$. Then we update $\theta_{A}$ with Eq. 5. After that, we fix $\theta_{A}$ and $\theta_{I}$ and update $\theta_{C}$ with Eq. 4 and repeat these steps until convergence.
@@
2:  Get the estimated sensitive attributes $\hat{si}$ with $f_{I}$
3:  Fix $\theta_{C}$ and $\theta_{A}$ and update $\theta_{I}$ using Eq. 6
4:  Update $\theta_{A}$ using Eq. 5
5:  Fix $\theta_{I}$ and $\theta_{A}$ and update $\theta_{C}$ using Eq. 4
@@
BFtS model consists of two GNNs and one DNN. If we consider GCN for training the GNN, the complexity is $\mathcal{O}(L|V|F^{2}+L|E|F)$, where $L$ is the number of layers and $F$ is the nodes in each layer (We assume that the number of nodes in each layer is the same) [4]. The complexity is polynomial, and therefore, BFtS is scalable. The complexity of FairGNN [14] is the same as the complexity of BFtS. FairGNN makes an additional forward pass to the trained sensitive attribute imputation network. Therefore, the empirical running time of FairGNN is higher than BFtS as described in the next section.
@@
Figure 7: Sensitivity analysis of $\alpha$ and $\beta$. BFtS is more sensitive to $\alpha$ than $\beta$. These parameters can be optimized using automated techniques, such as grid-search.
@@
Table 3: AVPR, F1 score, $\%\Delta DP$, and $\%\Delta EQOP$ of different methods. For the Bail, and NBA dataset, we outperform all baselines in terms of accuracy and fairness. For other datasets, we improve fairness while slightly sacrificing accuracy.
@@
Figure 9: 2-D kernel PCA node representations for the Bail dataset generated by the FairGNN and BFtS (our approach). We use different markers for nodes depending on their predicted class, sensitive attribute (race), and whether the sensitive attribute is missing. The results show that representations for nodes with missing values are more biased for FairGNN where they are concentrated in a negative class region. On the other hand, nodes with missing values are better spread over the space for BFtS representations. In BFtS, there are more "Race = Black" nodes than in FairGNN that have missing sensitive values in training but are predicted to be positive.
@@
Table 3 shows the average results for Bail, Credit, German, and NBA datasets. In terms of accuracy and fairness, BFtS outperforms all baseline approaches for the Bail and NBA datasets. We also outperform competitors on the Credit and German datasets in terms of fairness, with a slight AVPR loss. Although we lose some AVPR in the German and Credit datasets, we win in F1 and the AVPR loss is negligible when compared to the fairness benefit. Table 4 shows the average results for Bail, Credit, German, and NBA datasets. We have used GAT [67] to train $f_{C}$ and $f_{I}$. In terms of accuracy and fairness, BFtS outperforms all baseline approaches for the Bail dataset. We also outperform competitors on the Credit, German, and NBA datasets in terms of fairness with a slight AVPR loss. The AVPR loss is negligible when compared to the fairness benefit. Figure 9 shows kernel PCA node representations produced by FairGNN and BFtS using the Bail dataset. We show the samples with missing "Race" with a navy dot in the middle and samples predicted as positive class and negative class in green and red, respectively. The representations generated by FairGNN are noticeably more biased than the ones generated by our approach. A higher number of "Race = Black" samples with missing sensitive values in training are predicted as positive by BFtS than for FairGNN. Moreover, BFtS spreads nodes with missing values more uniformly over the space. This illustrates how the missing value imputation of FairGNN underestimates the bias in the training data and, therefore, the model could not overcome the bias in the predictions.
@@
We also vary the amount of observed data $\mathcal{V}_{S}$ and plot the results in Figure 8. We use $20\%,30\%,40\%$, and $50\%$ of training nodes as $\mathcal{V}_{S}$. For nearly all models, fairness increases and accuracy decreases with $|\mathcal{V}_{S}|$. With BFtS, both accuracy and bias often decline with the increase of $\mathcal{V}_{S}$. In the majority of the settings, our approach (BFtS) achieves a better fairness $\times$ accuracy trade-off than the baselines.
To see the impact of LDAM loss, we train $f_{I}$ with cross-entropy loss and compare the performance with $f_{I}$ trained with LDAM loss in Table 6. Evidently, LDAM achieves a better trade-off between accuracy and fairness.
Cross Entropy $f_{I}$ LDAM $f_{I}$
@@
Table 6: Ablation study using LDAM loss and cross-entropy loss for $f_{I}$. Using LDAM loss gives a better fairness and accuracy trade-off.
Ethical Considerations
Some of the ethical considerations related to our work include privacy, fairness, sustainability, and transparency. Our work is motivated by fairness and privacy. More specifically, missing sensitive values might be because users do not want to disclose their private data. Sustainability and transparency are not the focus of our work. However, we believe that existing methods for interpretable machine learning can improve the transparency of our model. The datasets used in our method are all public information and do not disclose private information from users.
Reproducibility
Our code can be found in an anonymous repository [^1] https://anonymous.4open.science/r/BFtS-6ADA.

### v5

@@ Appendix A Technical Appendices and Supplementary Material
$\min_{\theta_{class}}\sup_{\hat{s}}|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$
Let us assume that there can be $m$ different $\hat{s}$ that produce sensitive attributes $\hat{s}_{1},\hat{s}_{2},...,\hat{s}_{m}$. If we let $D_{i}$ be the uniform distribution over each $s_{i}$, then BFtS minimizes the $|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$ in the worst case over these different distributions $D_{i}$. Therefore, $f_{class}$ is a minimax estimator. The objective of Eq. 7 with optimal $\theta_{bias}^{*},\theta_{imp}^{*}$ and $f^{*}_{class}$ becomes:
@@
Therefore, $f_{class}$ is a minimax estimator and the maximal $\Delta DP$ of BFtS is minimum among all estimators of $s$.
At the global minima of $\theta_{class}$, $|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$ will be minimum. Its minimum value is 0. Therefore, $p(\mathbf{h}|\hat{s}=1)-p(\mathbf{h}|\hat{s}=0)=0$. The optimal $\theta^{*}_{class}$ achieves demographic parity for the worst case.
@@
If $s^{\prime}$ is a separate imputation independent of $f_{class}$ and $f_{bias}$, then the min max game according to Dai & Wang (2021) is the following:
$\begin{split}\min_{\theta_{class}}\max_{\theta_{bias}}~&\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|s^{\prime}=1)}[\log f_{bias}(\mathbf{h})]\\ &+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|s^{\prime}=0)}[\log(1-f_{bias}(\mathbf{h}))].\end{split}$
With optimal $\theta^{*}_{bias}$, the optimization problem simplifies to-
$\min_{\theta_{class}}-\log 4+JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0)).$
With independent imputation $s^{\prime}$, the adversary $f_{bias}$ tries to approximate the lower bound of Jensen Shannon Divergence (Weng, 2019). In BFtS, the adversarial imputation $f_{imp}$ tries to approximate the upper bound of the JS divergence, and the classifier $f_{class}$ tries to minimize the upper bound on the JS divergence provided by the adversarial imputation (see Proposition 1). Therefore,
@@
Let us consider a scenario when the independent missing sensitive value imputation (Dai & Wang, 2021) fails to converge. There are three possible reasons for the convergence to fail:
@@
In scenario $1$ and $2$ the supports of $p(\mathbf{h}|s^{\prime}=1)$ and $p(\mathbf{h}|s^{\prime}=0)$ are disjoint. According to (Dai & Wang, 2021), the optimization for the classifier with an optimal adversary is
$\min_{\theta_{class}}-\log 4+2JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$
@@
In scenario 3, $\hat{s}$ is uniform. Therefore, $p(\mathbf{h}|s^{\prime}=1)$ will be equal to $p(\mathbf{h}|s^{\prime}=0)$ which results in $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ being equal to 0. Moreover, the classifier may assign all training samples to a single class and yet have an adversary that may not be able to distinguish the sensitive attributes of the samples as $\hat{s}$ is random. If the classifier assigns all samples to the majority class, it will achieve low classification loss $\mathcal{L}_{class}$ along with minimum $\mathcal{L}_{bias}$. The training of the classifier may get stuck in this local minima (see Eq. 4). This phenomenon is similar to the mode collapse of GANs (Thanh-Tung & Tran, 2020).
@@
Algorithm 1 is a high-level description of the key steps applied for training BFtS. It receives the graph $\mathcal{G}$, labels $y$, labeled nodes $\mathcal{V}_{L}$, nodes with observed sensitive attributes $\mathcal{V}_{S}$ and hyperparameters, $\alpha$ and $\beta$ as inputs and outputs the GNN classifier $f_{class}$, sensitive attribute predictor, $f_{bias}$ and missing data imputation GNN $f_{imp}$. We first estimate $\hat{si}$ and then fix $\theta_{class}$ and $\theta_{bias}$ to update $\theta_{imp}$. Then we update $\theta_{bias}$ with Eq. 5. After that, we fix $\theta_{bias}$ and $\theta_{imp}$ and update $\theta_{class}$ with Eq. 4 and repeat these steps until convergence.
@@
2:  Get the estimated sensitive attributes $\hat{si}$ with $f_{imp}$
3:  Fix $\theta_{class}$ and $\theta_{bias}$ and update $\theta_{imp}$ using Eq. 6
4:  Update $\theta_{bias}$ using Eq. 5
5:  Fix $\theta_{imp}$ and $\theta_{bias}$ and update $\theta_{class}$ using Eq. 4
@@
BFtS model consists of two GNNs and one DNN. If we consider GCN for training the GNN, the time complexity is $\mathcal{O}(L|V|F^{2}+L|E|F)$, and space complexity is $\mathcal{O}(LE+LF^{2}+L|V|F)$, where $L$ is the number of layers and $F$ is the nodes in each layer (We assume that the number of nodes in each layer is the same) (Blakely et al., 2021). The complexity is polynomial, and therefore, BFtS is scalable. The complexity of FairGNN (Dai & Wang, 2021) is the same as the complexity of BFtS. FairGNN makes an additional forward pass to the trained sensitive attribute imputation network. Therefore, the empirical running time of FairGNN is higher than the one for BFtS, as shown in the next section.
@@
Figure 6 presents the convergence behavior of $\mathcal{L}_{class}$, $\mathcal{L}_{bias}$ and $\mathcal{L}_{imp}$ across training epochs on the NBA dataset.
Figure 6: Convergence of $\mathcal{L}_{class}$, $\mathcal{L}_{bias}$ and $\mathcal{L}_{imp}$. All three losses converge smoothly, confirming theoretical stability.
Figure 7: Sensitivity analysis of $\alpha$ and $\beta$. BFtS is more sensitive to $\alpha$ than $\beta$. These parameters can be optimized using automated techniques, such as grid search.
@@
$\%\Delta EQODDs$ ($\downarrow$) 11.56±0.02 6.15±0.03 4.81±0.04 3.04±0.02 4.51 ±2.13 3.49 ±0.05 2.05±0.04
Table 3: AVPR, F1 score, $\%\Delta DP$, % $\Delta$ EQOP, and % $\Delta$ EQODDs of different methods. For the Bail, and NBA dataset, we outperform all baselines in terms of accuracy and fairness. For other datasets, we improve fairness while slightly sacrificing accuracy.
@@
Figure 9: 2-D kernel PCA node representations for the Bail dataset generated by the FairGNN and BFtS (our approach). We use different markers for nodes depending on their predicted class, sensitive attribute (race), and whether the sensitive attribute is missing. The results show that representations for nodes with missing values are more biased for FairGNN where they are concentrated in a negative class region. On the other hand, nodes with missing values are better spread over the space for BFtS representations. In BFtS, there are more ‘Race = Black’ nodes than in FairGNN that have missing sensitive values in training but are predicted to be positive.
Worst-Case Fairness Accuracy Trade-off
As our proposed model operates under a worst-case fairness assumption, it may overestimate the bias in the complete data, as illustrated in Figure 2 using the Bail dataset. This results in a trade-off between fairness and utility, which is governed by a hyperparameter $\beta$. Figures 10(a), (b), and (c) show the F1 score, $1-\Delta\text{DP}$, and $1-\Delta\text{EQOP}$, respectively, as $\beta$ varies and with $30\%$ of sensitive values observed. Here, F1 serves as the utility metric, while $1-\Delta\text{DP}$ and $1-\Delta\text{EQOP}$ quantify fairness. We compare our BFtS model against a fair adversarial model trained with complete sensitive information. By tuning $\beta$, we balance predictive utility and fairness under worst-case bias, achieving an effective trade-off.
@@
Figure 11: Fairness accuracy trade off for large scale graph dataset. BFtS achieves better fairness-accuracy trade-off
[…]
BFtS 0.71 $\pm$ 0.10 0.15 $\pm$ 0.02
Vanilla 0.65 $\pm$ 0.10 0.21 $\pm$ 0.01
Table 6: Fair regression: MSE and Fairness ($\Delta DP$) Comparison Between BFtS and Vanilla Models
Large Scale Graph Dataset
To verify the performance of BFtS on large-scale dataset, we generated a synthetic graph with 250k nodes and 100 features, as we were unable to identify a large-scale real-world graph dataset with sensitive values. Following the setup in Figure 4 (with 150k nodes in the non-protected group and assortativity = 0.77). We compared BFtS against other baselines in Figure 11 across different hyperparameter settings. BFtS completed training in 85.4 minutes versus RNF’s (the strongest baseline in this scenario) 114.6 minutes.
Average Performance with GCN and GAT
Table 3 shows the average results for Bail, Credit, German, and NBA datasets. In terms of accuracy and fairness, BFtS outperforms all baseline approaches for the Bail and NBA datasets. We also outperform competitors on the Credit and German datasets in terms of fairness, with a slight AVPR loss. Although we lose some AVPR in the German and Credit datasets, we win in F1 and the AVPR loss is negligible when compared to the fairness benefit. Table 4 shows the average results for Bail, Credit, German, and NBA datasets using GAT Velickovic et al. (2018) to train $f_{class}$ and $f_{imp}$. In terms of accuracy and fairness, BFtS outperforms all baseline approaches for the Bail dataset. We also outperform competitors on the Credit, German, and NBA datasets in terms of fairness with a slight AVPR loss. The AVPR loss is negligible when compared to the fairness benefit.
Visualization of Representations
Figure 9 shows kernel PCA node representations produced by FairGNN and BFtS using the Bail dataset. We show the samples with missing ”Race” with a navy dot in the middle and samples predicted as positive class and negative class in green and red, respectively. The representations generated by FairGNN are noticeably more biased than the ones generated by our approach. A higher number of ”Race = Black” samples with missing sensitive values in training are predicted as positive by BFtS than for FairGNN. Moreover, BFtS spreads nodes with missing values more uniformly over the space. This illustrates how the missing value imputation of FairGNN underestimates the bias in the training data and, therefore, the model could not overcome the bias in the predictions.
Fair node regression
While fairness-aware node regression datasets are lacking, we created a synthetic node regression task (assortativity = 0.77) following our setup in Figure 5, with group-wise targets from $\mathcal{N}(0,1)$ and $\mathcal{N}(2,1)$. The Table 6 reports MSE and $\Delta DP$ (Berk et al., 2017) for regression.
@@
We also vary the amount of observed data $\mathcal{V}_{S}$ and plot the results in Figure 8. We use $10\%,20\%,...,80\%$, of training nodes as $\mathcal{V}_{S}$. For nearly all models, fairness increases and accuracy decreases with $|\mathcal{V}_{S}|$. With BFtS, both accuracy and bias often decline with the increase of $\mathcal{V}_{S}$. In the majority of the settings, our approach (BFtS) achieves a better fairness $\times$ accuracy trade-off than the baselines.
To see the impact of LDAM loss, we train $f_{imp}$ with cross-entropy loss and compare the performance with $f_{imp}$ trained with LDAM loss in Table 7. Evidently, LDAM achieves a better trade-off between accuracy and fairness.
Cross Entropy $f_{imp}$ LDAM $f_{imp}$
@@
Table 7: Ablation study using LDAM loss and cross-entropy loss for $f_{imp}$. Using LDAM loss gives a better fairness and accuracy trade-off.